Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2024 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn  Not logged in ELOG logo
Message ID: 171     Entry time: 16 Sep 2024, 02:21
Author: Wonhyuk Jo 
Group: MID 
Subject: Shift summary 

- Pump-probe experiments are being conducted using PuMa and PP laser. 
- The user program is ongoing over the night.

Tomorrow, 
- Rotating the sample phi angle. The sample holder will be retracted from the MPC.
- Search the asymmetric film peak and conduct the PuMa and PP laser experiment.

ELOG V3.1.4-7c3fd00